### v5

@@ 6 Empirical Study
In this empirical study, we apply our theoretical results to real-world trade data spanning from 2000 to 2019. This analysis provides quantitative evidence to support our previous arguments, making the theoretical insights more tangible and relevant for policymakers. The empirical analysis reveals trends in national competitiveness and bargaining power over time, highlighting how different countries have gained or lost. This provides a historical perspective on the impacts of globalization and trade wars.
Between 2000 and 2019, the world economy underwent significant transformations and challenges, albeit with varying degrees of success across different regions. This period witnessed a substantial increase in global trade volumes, driven by rapid technological advancements, the expansion of international trade agreements, and the rise of emerging markets. The integration of China into the WTO in 2001 was a pivotal moment, leading to a surge in global trade activities. The rise of emerging markets, particularly China and India, played a crucial role in global economic growth but also led to a trade war between China and the USA. The 2008 financial crisis and the subsequent European debt crisis were major turning points, causing widespread economic downturns and leading to substantial policy interventions by governments and central banks worldwide.
As trade volumes increased with globalization, nations demanded more of the world’s reserve currency, primarily through trade surpluses. Consequently, the reserve currency appreciates, decreasing the currency issuer’s exports and increasing trade deficits. The issuer may have three options: encouraging local currency settlement in international transactions or reducing its currency’s usage as the world’s reserve currency; ignoring the trade deficits and focusing on national competitiveness while still increasing exports; or rolling back globalization worldwide. For the USA, reshoring its manufacturing sectors would enhance its competitiveness and grow exports; however, trade deficit may still be an issue according to the impossibility trilemma. For China, significant growth is necessary to shrink the competitiveness gap with the USA, according to the Matthew effect in $\pi$.
@@
This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$. The bilateral trade data are sourced from the United Nations ComTrade database (2021). We use import and export data for non-consumer goods from the World Bank’s WITS Database (2021). For comparison, we focus on the years 2000, before China entered the WTO in 2001, and 2017, before the beginning of the China-USA trade war in 2018. Though all economies in the data are used in the calculations, we only present the results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the UK, and the USA). Their ISO-3 country codes are CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA, respectively. We use $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ in the computations. The results are in the following tables and figures.
Figure IV: Competitiveness $\pi_{i}$ for China, Russia, and G7 countries (2000-2019).
@@
Figure IV illustrates the competitiveness trends from 2000 to 2019. During this period, China and Russia experienced significant gains, while Canada remained relatively stable, and the other six countries saw declines. Notably, the USA’s competitiveness grew steadily until 2003, followed by a sharp decline until 2011, and a gradual recovery thereafter, resulting in an overall decrease of 12.7%. China’s competitiveness surged during the two crises but plateaued otherwise. Canada and Russia demonstrated resilience due to rising commodity prices during transformative years. The four Western European countries were significantly impacted by the debt crisis, with their competitiveness continuing to fall until 2015. Germany, focused on manufacturing, began its decline in 2000, while the other three countries started in 2008. Japan’s competitiveness nearly halved in the first eight years but stabilized between 2007 and 2014, suffering a domino effect from 2000 to 2007 as China and South Korea challenged its automobile and electric industries. Beyond these nine countries, others collectively improved their competitiveness by 22.7%, as depicted in Figure VI. Many Southern Asian, Southeast Asian, and Middle East countries, including India, Indonesia, Iraq, United Arab Emirates, and Vietnam, tripled their competitiveness.
@@
The Matthew effect is clearly demonstrated. For instance, the USA and China’s production ratio was $1.344$ in 2018, but their competitiveness ratio was $3.070$. This effect is further evidenced by Figure VI, which shows a strong positive correlation between $\pi_{i}$ and $g_{i}^{2}$ after controlling $g_{i}$. The partial correlation, however, declined from .757.757.757.757 to .527.527.527.527 during the study period when the trade system shifted from less to more multipolar.
Figure V: The sum of $\pi_{i}$ for all other countries (2000-2019).
Figure VI: Correlation between $\pi_{i}$ and $g^{2}_{i}$ after controlling $g_{i}$ (2000-2019).
@@
* The numerators are for 2019 and the denominators for 2017.
A straightforward implication of Eq. (4) is country $j$’s direct contribution to country $i$’s competitiveness in percentage, i.e., $\pi_{j}P_{ji}/\pi_{i}\times 100\%$. Table I lists these contributions for 2017 and 2019. According to the diagonal numbers, the USA was the most self-sufficient in feeding its competitiveness, with a slight increase from 2017 to 2019. The ninth column also shows that the USA was the most significant contributor to Canada, China, Germany, Japan, and the UK in 2019. For the same year, China was the most significant contributor to Russia’s competitiveness and the second largest to those of Canada, Japan, and the USA. France, Germany, and Italy were highly reliant on each other’s contributions, and their average dependence on Russia was 8.7 times higher than the USA’s dependence on Russia. In value, the USA contributed .14492.025965=.003763.14492.025965.003763.14492*.025965=.003763.14492 ∗ .025965 = .003763 to Canada while Canada contributed .013878.293880=.004078.013878.293880.004078.013878*.293880=.004078.013878 ∗ .293880 = .004078 to the USA in 2019. Comparing the years before and after the beginning of the China-USA trade war, the USA’s contributions to China dropped by $21.7\%$, calculated as 1(1.5849.093905)/(2.1950.086596)11.5849.0939052.1950.0865961-(1.5849*.093905)/(2.1950*.086596)1 - ( 1.5849 ∗ .093905 ) / ( 2.1950 ∗ .086596 ), while China’s contribution to the USA reduced by $15.2\%$, calculated as 1(.8107.29388)/(1.0322.272142)1.8107.293881.0322.2721421-(.8107*.29388)/(1.0322*.272142)1 - ( .8107 ∗ .29388 ) / ( 1.0322 ∗ .272142 ).
Figure VII: Bargaining power $\beta_{i}$ (2000-2019).
Figure VII illustrates the national bargaining power from 2000 to 2019, displaying patterns similar to those in Figure IV, albeit smoother. China’s bargaining power steadily increased due to rapid production growth, even during its $\pi_{i}$’s troughs in 2005 and 2015. Notably, China’s bargaining power consistently surpassed its competitiveness, whereas the USA’s bargaining power was consistently lower than its competitiveness, indicating $\beta$’s mitigation of the Matthew effect in $\pi$. Russia’s bargaining power sharply declined following the Russo-Georgian War in 2008 and the annexation of Crimea in 2014.
Figure VIII illustrates the comparative bargaining power of the USA against other countries. The bargaining power of four Western European countries experienced a slight increase following the establishment of the Eurozone in 1999, but saw a modest decline post-2008, culminating in an average loss of 15.7% of their bargaining power to the USA. Conversely, China and Russia significantly enhanced their relative power multiple times within the first 15 years, maintaining a steady level thereafter. Japan, however, suffered a substantial reduction, losing approximately 50% of its relative bargaining power. Canada’s relative bargaining strength surged during crises but swiftly reverted to its original level, attributable to its strong alignment with the USA.
Figure VIII: Bargaining power versus the USA (2000-2019).
@@
Table II presents the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for the years 2000 and 2017, calculated from Eqs. (9) and (12). For instance, a 1% reduction in imports from China by the USA in 2017 would have resulted in an increase of .04889% in the USA’s competitiveness. Although the USA may have had reasons to initiate trade wars against all eight other countries, its primary targets were China in 2017 and Japan in 2000, as indicated by the numerators and denominators in the last row. As evidenced by the last column, China and Japan would have been the two largest beneficiaries had they further collaborated with the USA in 2017 and 2000, respectively. However, a comparison of the numerators and denominators in the column reveals a considerable reduction in benefits from 2000 to 2017. Additionally, initiating trade friction from any of these eight countries against the USA would result in a loss for the initiating country. A trade war against Russia would be nearly fruitless due to the small numerators and denominators in the eighth column. Furthermore, the numbers symmetric with the diagonal exhibit opposite signs, corroborating Corollary 3.
@@
Table III lists the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated from Eqs. (17), (18), and (12). From the diagonal, globalization significantly benefited China, Germany, Italy, Japan, and Russia. It was advantageous for Canada in 2000 but detrimental in 2017. The effects on France were insignificant due to small numerator and denominator. The USA and the UK would be anti-globalization advocates, explaining recent events such as the Trump administration’s withdrawal from international organizations and the UK’s Brexit. Though business interest groups overwhelmingly profited from the gigantic international markets and low-priced raw materials, growing low-income Americans from manufacturing offshoring demanded more inexpensive goods from China, uplifting trade deficits and national debts but benefiting from low inflation and affordable merchandises. Also, China was the largest beneficiary of globalization in 2017 with a 2.324% hike in $\pi_{i}$ for a 1% decrease in $P_{ii}$, so its Belt and Road Initiative (BRI) would have a positive impulse on its $\pi_{i}$. Russia was the largest beneficiary of globalization in 2000.
@@
The side effects on third parties are positioned off the diagonal. For instance, the USA’s further globalization benefited all eight other countries, thereby making its protectionist policies likely to face significant resistance from its partners. China emerged as the largest beneficiary of the USA’s globalization in both years. Conversely, Canada experienced the least impact from the USA’s protectionism in 2017; its further globalization would have marginally benefited the USA, as indicated by the first row. The last column suggests that further globalization from seven other countries would adversely affect the USA; Japan and China topped the list in 2000 and 2017, respectively. Their adverse impact decreased by 70.2% and increased by 380%, respectively, over the 17 years. Additionally, the second row elucidates the USA and its Western allies’ objection to China’s BRI-like projects, which intensified by $4.74$ times from 2000 to 2017, according to the ratios of their numerators and denominators.
@@
We assess the externalities on the third parties due to the USA’s trade war against China in 2018 and 2019. Table IV is derived from Eqs. (9), (11), and (12). When the USA reduced its imports from China by 1% in 2018 and 2019, China’s $\pi_{j}$ suffered a significant loss of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ earned a non-significant gain of .0540% and .0460%, respectively. Canada was a free rider, and the other six countries were slightly negatively affected in 2018 and 2019. Consequently, these countries would likely refrain from taking formal action to prevent the trade war. Furthermore, Canada or Russia would be the best ally with the USA or China, respectively, in this conflict. The second-best allies would be the UK for the USA and Japan for China, respectively. From 2018 to 2019, the effects were slightly mitigated for all these countries. Likely recognizing the mutually destructive consequences of the trade war, China and the USA agreed on a new trade deal in early 2020 (cf, US Trade Representative, 2020), concluding the first phase of a long-term conflict. However, the cumulative effects would be significant as the conflict is expected to continue for years.
@@
Table IV: $1000\times\frac{\mathrm{d}\log\pi_{k}}{\mathrm{d}\log P_{ji}}$ for the China-USA Trade War’s side effects on third parties
@@
* Country $i$ is for the USA, $j$ for China, and $k$ for others.
Finally, the cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of this conflict in 2018,
@@
For a positive $\mathrm{d}P_{ji}$, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V). Thus, an economic conflict with a negative $\mathrm{d}P_{ji}$ was possible and, in this lose-lose conflict, the Nash bargaining solution was:
[…]
@@
for 2018. For each dollar’s decrease in exports from China to the USA, $\mathrm{d}P_{ji}=-\$1/g_{j}$. According to the bargaining solution, China would have changed its imports from the USA by $\mathrm{d}P_{ij}=\lambda_{ji}^{*}\mathrm{d}P_{ji}$ of the USA’s production, which is .$66164.49233*(-1/g_{j})*g_{i}=-\$.$66164. Therefore, if this resolution had been applied, the USA would have earned \$.33836=(-.66164)-(-1) surplus for each decreased dollar of imports from China. Additionally, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}<0$. To maintain its competitive advantage $\pi_{i}$ and reduce national debt by earning a trade surplus in the bilateral friction, by the impossibility trilemma in Theorem 4 , the USA had to sacrifice its exports and shrink its imports — in either case, $\mathrm{d}P_{ij}<0$.
7 Discussions
Economic globalization and trade wars have emerged as significant public interests due to global competition among nations, large-scale restructuring of labor markets, and the international distribution of added value. This study posits that import and export data already reflect these issues, and that competitive advantages, rather than comparative advantages, are the primary drivers of trade wars and de-globalization. Consequently, we aim to extract a nation’s competitive strength and bargaining power from the data, identify optimal strategies for addressing trade conflicts and cooperation, and determine the extent to which a nation should pursue globalization.

### v6

@@ 6 Empirical Study
In this empirical study, we apply our theoretical framework to real-world trade data spanning from 2000 to 2019. This analysis offers quantitative evidence that supports our earlier arguments, making the theoretical insights more concrete and relevant for policymakers. The empirical findings reveals evolving trends in national competitiveness and bargaining power, illustrating how various countries have gained or lost ground over time. These insights provide a historical lens through which to examine the impacts of globalization and trade conflicts.
Between 2000 and 2019, the global economy experienced profound transformations and challenges, with varying outcomes across regions. This period saw a marked increase in global trade volumes, fueled by rapid technological advancements, the expansion of international trade agreements, and the emergence of new economic powers. A pivotal moment was China’s accession to the WTO in 2001, which triggered a surge in global trade activity.
The rise of emerging markets — particularly China and India — was instrumental in driving global economic growth. However, it also contributed to escalating trade tensions, most notably between China and the USA. The 2008 financial crisis and the subsequent European debt crisis were major inflection points, causing widespread economic downturns and prompting significant policy interventions by governments and central banks worldwide.
As globalization intensified and trade volumes expanded, nations increasingly demanded the world’s reserve currency, primarily through trade surpluses. This demand led to appreciation of the reserve currency, which in turn reduced the currency issuer’s export competitiveness and widened trade deficits. Thus, reversing globalization trends could be a viable option for the issuer. For the USA, reshoring manufacturing sectors could reduce imports and deficits. For China, sustained economic growth is essential to narrow the competitiveness gap with the USA, in line with the Matthew effect in $\pi$.
@@
This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$. Bilateral trade data are sourced from the United Nations ComTrade database (2021), while import and export data for non-consumer goods come from the World Bank’s WITS Database (2021). For comparison, we focus on two benchmark years: 2000, prior to China’s accession to the WTO in 2001, and 2017, just before the onset of the China-U.S. trade war in 2018.
Although the calculations include all economies in the dataset, we present results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the UK, and the USA), identified by their ISO-3 codes: CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA. The computations use the default coefficients, $\lambda_{ji}=\lambda_{ji}^{*}$, and the results are summarized in the following tables and figures.
Figure IV: Competitiveness $\pi_{i}$ for China, Russia, and G7 countries (2000-2019)
@@
Figure IV illustrates competitiveness trends from 2000 to 2019. During this period, China and Russia recorded significant gains, while Canada remained relatively stable, and the other six countries experienced declines. Notably, the USA’s competitiveness rose steadily until 2003, then dropped sharply through 2011, followed by a gradual recovery — resulting in an overall decline of 12.7%. China’s competitiveness surged during the two major crises but otherwise plateaued. Canada and Russia demonstrated resilience, supported by rising commodity prices. The four Western European countries were severely affected by the debt crisis, with competitiveness falling until 2015. Germany’s decline began in 2000, while the other three started in 2008. Japan’s competitiveness nearly halved in the first eight years, then stabilized between 2007 and 2014. From 2000 to 2007, Japan faced a reverse Matthew effect as China and South Korea challenged its automobile and electric industries.
Beyond these nine countries, other economies collectively improved competitiveness by 22.7%, as shown in Figure VI. Many countries in Southern Asian, Southeast Asian, and the Middle East — including India, Indonesia, Iraq, the United Arab Emirates, and Vietnam — tripled their competitiveness.
@@
The Matthew effect is clearly evident in the data. For example, in 2018, the production ratio between the USA and China was $1.344$, but their competitiveness ratio reached $3.070$. This disparity underscores the cumulative advantage phenomenon. Further evidence in presented in Figure VI, which illustrates a strong positive correlation between $\pi_{i}$ and $g_{i}^{2}$, after controlling for $g_{i}$. However, the partial correlation declined from .757.757 to .527.527 over the study period, coinciding with the transition of the trade system from a less multipolar to a more multipolar structure.
Figure V: Sum of $\pi_{i}$ for all other countries (2000-2019)
Figure VI: Correlation between $\pi_{i}$ and $g^{2}_{i}$ after controlling for $g_{i}$ (2000-2019)
@@
* The numerators correspond to 2019 values, and the denominators to 2017.
A direct implication of Eq. (4) is that country $j$’s contribution to country $i$’s competitiveness can be expressed as a percentage: $\pi_{j}P_{ji}/\pi_{i}\times 100\%$. Table I reports these contributions for 2017 and 2019. The diagonal values indicate that the USA was the most self-sufficient in sustaining its competitiveness, with a slight increase from 2017 to 2019. The ninth column further shows that, in 2019, the USA was the largest contributor to Canada, China, Germany, Japan, and the UK. In the same year, China was the primary contributor to Russia’s competitiveness and the second-largest contributor to Canada, Japan, and the USA. France, Germany, and Italy exhibited strong mutual reliance, while their average dependence on Russia was 8.7 times higher than the USA’s dependence on Russia.
Bilateral break-even points in Eq. (6), in general, do not exist. For example, in 2019, the USA contributed .14492.025965=.003763.14492*.025965=.003763 to Canada, whereas Canada contributed .013878.293880=.004078.013878*.293880=.004078 to the USA. Comparing the periods before and after the onset of the China-USA trade war, the USA’s contributions to China decreased by $21.7\%$, calculated as 1(1.5849.093905)/(2.1950.086596)1-(1.5849*.093905)/(2.1950*.086596), while China’s contribution to the USA fell by $15.2\%$, calculated as 1(.8107.29388)/(1.0322.272142)1-(.8107*.29388)/(1.0322*.272142).
Figure VII: Bargaining power $\beta_{i}$ (2000-2019)
Figure VII illustrates national bargaining power from 2000 to 2019, showing patterns similar to Figure IV but smoother. China’s bargaining power rose steadily due to rapid production growth, even during troughs in its $\pi_{i}$ in 2005 and 2015. Notably, China’s bargaining power consistently exceeded its competitiveness, whereas the USA’s bargaining power was consistently lower than its competitiveness — evidence of $\beta$’s role in mitigating the Matthew effect in $\pi$. Russia’s bargaining power declined sharply after the Russo-Georgian War in 2008 and the annexation of Crimea in 2014.
Figure VIII compares the USA’s bargaining power against the other countries. Western European countries saw a slight increase after the Eurozone’s formation in 1999, followed by a modest decline post-2008, resulting in an average 15.7% loss of bargaining power relative to the USA. Conversely, China and Russia significantly strengthened their relative power during the first 15 years, then stabilized. Japan, however, experienced a substantial decline, losing about 50% of its relative bargaining power. Canada’s bargaining strength spiked during crises but quickly reverted to its original level, reflecting its strong alignment with the USA.
Figure VIII: Bargaining power versus the USA (2000-2019)
@@
Table II reports the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for the years 2000 and 2017, calculated using Eqs. (9) and (12). For example, a 1% reduction in U.S. imports from China in 2017 would have increased U.S. competitiveness by .04889%. Although the USA may have had reasons to initiate trade disputes with all eight other countries, its primary targets were China in 2017 and Japan in 2000, as indicated by the numerators and denominators in the last row. The last column shows that China and Japan would have been the two largest beneficiaries had they strengthened collaboration with the U.S. in 2017 and 2000, respectively. However, comparing the numerators and denominators across years reveals a significant decline in potential benefits from 2000 to 2017. Moreover, any of these eight countries initiating trade friction against the USA would incur losses. A trade war with Russia, in particular, would be almost fruitless due to the small values in the eighth column. Finally, the numbers symmetric to the diagonal exhibit opposite signs, confirming Corollary 3.
@@
Table III presents the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated using Eqs. (17), (18), and (12). The diagonal values indicates that globalization significantly benefited China, Germany, Italy, Japan, and Russia. Canada gained in 2000 but faced negative effects in 2017. For France, the impact was negligible due to small numerator and denominator values. The USA and the UK appeared as anti-globalization advocates, which aligned with recent developments such as the Trump administration’s withdrawal from international organizations and the UK’s Brexit. While business interest groups profited from vast international markets and low-cost raw materials, many low-income Americans — affected by manufacturing offshoring — demanded cheaper goods from China. This dynamic increased trade deficits and national debts but also delivered low inflation and affordable consumer products. China emerged as the largest beneficiary of globalization in 2017, with a 2.324% increase in $\pi_{i}$ for every 1% decrease in $P_{ii}$. This advantage underscores the positive momentum behind its Belt and Road Initiative (BRI). In contrast, Russia was the largest winner in 2000.
@@
Off-diagonal values reveal spillover effects on third parties. For example, further globalization by the USA benefited all eight other countries, suggesting that U.S. protectionist policies would likely face resistance from its partners. China consistently gained the most from U.S. globalization in both years. Conversely, Canada experienced minimal impact from U.S. protectionism in 2017, and its own globalization would have only marginally benefited the USA, as indicated in the first row. The last column shows that additional globalization from seven other countries would negatively affect the USA, with Japan and China exerting the greatest adverse impact in 2000 and 2017, respectively. Over 17 years, Japan’s negative effect decreased by 70.2%, while China’s increased by 380%. Furthermore, the second row highlights the growing opposition of the USA and its Western allies to China’s BRI-like projects, which intensified by a factor of $4.74$ between 2000 and 2017.
@@
We assess the externalities imposed on third parties by the China-USA trade war during 2018 and 2019. Table IV is derived from Eqs. (9), (11), and (12). When the USA reduced its imports from China by 1% in 2018 and 2019, China’s competitiveness $\pi_{j}$ suffered significant losses of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ experienced non-significant gains of .0540% and .0460%, respectively. Canada benefited as a free rider, whereas the other six countries were slightly negatively affected in both years. From 2018 to 2019, the negative effects were slightly mitigated for all countries. Consequently, these countries were unlikely to take immediate action to prevent the dispute. However, cumulative spillover effects are expected to be significant as the conflict is likely to persist for years. Among potential allies, Canada and Russia would have been the best partners for the USA and China, respectively. The second-best allies would have been the UK for the USA and Japan for China.
@@
Table IV: $1000\times\frac{\mathrm{d}\log\pi_{k}}{\mathrm{d}\log P_{ji}}$ — Side Effects of the China-USA Trade War on Third Parties
@@
* Country $i$ refers to the USA, $j$ to China, and $k$ to others.
Finally, the cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of the conflict in 2018:
@@
For a positive $\Delta P_{ji}$, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V). Thus, an economic conflict with a negative $\Delta P_{ji}$ was possible, and in this lose-lose scenario, the Nash bargaining solution was:
$\lambda_{ji}^{*}=\sqrt{\frac{.092618\times 17494.79}{.284327\times 23511.17}}=.49233.$
@@
For each dollar decrease in exports from China to the USA ($\Delta P_{ji}=-\mathdollar 1/g_{j}$), China would have adjusted its imports from the USA by: $\Delta P_{ij}=\lambda_{ji}^{*}\Delta P_{ji}$ of the USA’s production, which is $49233*(-1/g_{j})*g_{i}=-\mathdollar.66164.$ Therefore, if this resolution had been applied, the USA would have earned a surplus of .33836=(.66164)(1).33836=(-.66164)-(-1) for each dollar of reduced imports from China. Additionally, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\Delta P_{ji}<0$.
To maintain competitive advantage and reduce national debt by earning trade surpluses, the USA, by the impossibility trilemma, had to reduce its imports and exports — i.e., $\Delta P_{ji}<0$ — through policies such as tariffs and manufacturing reshoring.
7 Policy Implications and Discussions
Economic globalization and trade wars have become pressing public concerns, driven by intensifying global competition, large-scale labor market restructuring, and the international distribution of added value. This study argues that import and export data already capture these dynamics and that competitive advantage — rather than comparative advantage — is the primary force behind trade wars and the trend toward deglobalization.
Our objective is to extract insights into a nation’s competitive strength and bargaining power from trade data, identify optimal strategies for managing trade conflicts and fostering cooperation, and assess the appropriate level of engagement in globalization. When applying this policy-oriented framework in real-world trade scenarios, it is essential to explore alternative approaches that adapt and extend the framework to specific national and economic contexts.
7.1 Practical and Policy Implications
This paper conceptualizes national competitiveness as a systemic property, rather than a simple reflection of unilateral production metrics or bilateral trade balances. Specifically, the metric $\pi$ represents stationary standings among nations, derived from their economic interactions — namely, bilateral trade flows. This formulation positions $\pi$ as a network-wide measure of economic strength, accounting for both direct and indirect effects across global value chains. Policymakers should monitor $\pi$ and its fluctuations alongside trade balances and production metrics to gain a more comprehensive understanding of national competitiveness.
7.1.1 Prioritizing Trade Objectives
When navigating the trade-off between competitiveness and trade balance, two actionable ratios can guide bilateral trade decisions: long-run competitiveness-neutral point $\lambda_{ji}=\dfrac{\pi_{j}}{\pi_{i}}$, useful for managing strategic rivalries without escalation; short-run trade-balance-neutral point $\lambda_{ji}=\dfrac{g_{j}}{g_{i}}$, suitable for addressing immediate concerns such as employment, debt sustainability, or foreign exchange reserves. These benchmarks help trade negotiators align policy actions with either long-term strategic goals or short-term economic stability.
The choice between strategies based on competitive versus comparative advantage should reflect national characteristics. Small economies are advised to adopt comparative advantage strategies to promote inclusive growth and mutual benefit, while superpowers may pursue competitive advantage strategies that favor dominant players. Most countries will adopt mixed, country-specific objectives. For example, in 2020, the USA faced strategic challenges from China alongside massive national debt.
Yet, competitiveness is often overlooked by incumbent administrations due to its intangible nature and long-term impact. Few quantitative metrics like $\pi$ have been used to measure competitiveness growth.
7.1.2 Recognizing Asymmetries Early
The convex nature of competitiveness creates reinforcing dynamics — often referred to as the Matthew effect — which makes it difficult for latecomers to catch up unless they grow substantially faster. This has critical implications for strategic sequencing, such as determining when to confront or collaborate with emerging rivals. Convexity also increases the risk of a Thucydides trap, where instability arises from competition between the two leading superpowers. Stability is more likely when a single superpower coexists with several great powers. Our empirical studies suggest that such instability may persist due to ongoing competition between China and the USA.
Consequently, incumbent superpowers should prioritize maintaining their leadership position. Meanwhile, small, open economies can leverage agglomeration effects by aligning with high-$\pi_{i}$ hubs and using pricing strategies to offset biases toward large markets. Fast-growing challengers aiming to close the competitiveness gap should avoid systematic confrontations with significantly larger economies.
Another critical asymmetry is the impossibility trilemma, which clarifies trade-offs often obscured in policy debates. Large economies should use it to set realistic bilateral goals, recognizing that achieving both a surplus and higher competitiveness may require reducing exports to challengers or broadly curbing imports.
7.1.3 Competition and Cooperation Corridors
The decision to compete or collaborate with a specific country depends on the marginal impact of imports and exports on $\pi_{i}$. Given limited resources, country $i$ may prioritize its top competitors and collaborators while ignoring others.
The counterbalance equilibrium provides a formal rationale for mixed trade policies that combine competitive pressure with cooperative incentives. This unity of opposites suggests avoiding extreme actions in both retaliations and cooperation. For example, controlling rare minerals and advanced technologies could escalate issues to national security levels.
Trade relationships evolve over time, making timing critical. Countries should develop an early-warning dashboard that annually tracks changes in $\pi$, $\beta$, and derivatives in Theorems 2 and 6. Thresholds — such as elasticities exceeding $\pm.1\%$ — should trigger interagency reviews if signs flip or magnitudes shift significantly.
7.1.4 Negotiation leverage and institutional design
The square-root formulation in $\lambda_{ji}^{*}$ symmetrically balances competitiveness and trade balance. It is scale-free and can serve as a default “fair retaliation” coefficient in bilateral trade protocols. However, this solution can be strategically manipulated (Corollary 6), particularly if a weaker nation appears to avoid competition with a dominant one, which may overlook this trap. The asymmetric Nash solution in Eq. (16) also benefits smaller countries more than the non-cooperative solutions in Corollary 5.
Moreover, when analyzing the relationship between a weaker country $j$ and a dominant country $i$, the Matthew effect indicates that the Nash bargaining solution yields $\lambda_{ji}^{*}>\pi_{j}/\pi_{i}$. Consequently, the United States may identify only a limited number of countries as suitable partners for deeper collaboration in its pursuit of enhanced national competitiveness (see Section 3.2). Nevertheless, all other economies can benefit from cooperating with the sole superpower, as outlined in Corollary 1 and supported by the empirical study.
Both the bilateral solution $\lambda_{ji}^{*}$ and the multilateral solution $\beta_{i}$ faces drawbacks due to simplifying assumptions that often fail to capture the complexity of international relations. Incorporating unequal bargaining power and third-party influence can improve these solutions. For instance, the USA may impose stricter retaliation measures (e.g., tariffs and quotas) on nations running trade surpluses with it. This may exclude China, given their comparable production sizes, with or without purchasing power parity.
7.1.5 Globalization and Supply-Chain strategy
Before adjusting border measures, country $i$ should assess its sensitivity to changes in its domestic share $P_{ii}$, i.e., $\dfrac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$. This indicates whether further globalization or protectionism enhances competitiveness. This empirical approach offers a pragmatic alternative to ideological debates over liberation versus reshoring.
A mixed strategy appears optimal: deepen ties with partners in $\mathcal{N}_{i}^{+}$ and selectively ring-fence against those in $\mathcal{N}_{i}^{-}$. In practice, however, both sets are too large. Accordingly, country $i$ should consider joining or upgrading PTAs/RTAs rich in $\mathcal{N}_{i}^{+}$, while negotiating carve-outs for areas with high $\mathcal{N}_{i}^{-}$ exposure — such as rules-of-origin thresholds or snap-back clauses. With these agreements, the spillover effects in Eqs. (11) and (18) require adjustments.
Due to the 2018 China-USA trade war, fragmentation between two blocs of nations has increased tension in the international trade system. The first bloc includes the over-globalized USA and its traditional allies, while the second comprises nations heavily reliant on China’s infrastructure-building capacity and affordable goods. Trade between blocs may decrease, while intra-bloc trade may increase. Countries outside these blocs may benefit.
7.1.6 Third-party spillovers and Trade Alliances
Partners play a pivotal role in the trade network. While uniform trade policies rarely benefit all partners, heterogeneous policies can generate gains across countries. Several methods — such as Eqs. (6) and (12), or the sets $\mathcal{N}^{+}_{i}$ and $\mathcal{N}^{-}_{i}$ — help identify competitors and collaborators.
Given any partner, country $i$ can enhance its competitiveness by adopting targeted strategies: weakening ties with competitors or strengthening ties with collaborators. These strategies adjust existing trade relationships, and before implementing them, a country should anticipate who stands to gain and who might resist — critical for coalition-building and mitigating opposition.
Gains and losses extend beyond principal actors to third-party countries. To manage these dynamics, countries should use Eqs. (11) and (18) to measure spillover effects, forecast ally reactions, and design compensatory mechanisms — thereby reducing coalition leakage during disputes. Ignoring such side effects can turn collaborators into competitors. Partner-specific collaboration can yield substantial returns, but these benefits often diminish over time, underscoring the need to regularly update partner rankings and reassess traditional alliances using current data.
7.2 Discussions and Extensions